# High Dimensional Ensemble Kalman Filter

### v1

@@ 8 Discussion
In this paper, we propose high dimensional EnKF algorithms to suit the needs of high resolution data assimilation. We take the opportunity to study the theoretical properties of the EnKF under both the fixed and the high dimensional Gaussian state-space model (2.8). The proposed high dimensional EnKF methods and algorithms can provide consistent data assimilation results in the high dimensional setting. The investigation of the high dimensional EnKF when the physical model $\mathcal{M}_{t}$ is misspecified gives a description of the performance of the EnKF under misspecified models.
The study shows that the proposed high dimensional EnKF can replace the widely used inflation method for data assimilation of a large dynamic system based on two considerations. One is that the established theoretical analysis on the consistency and the error bounds makes the proposed high dimensional methods more understandable with a theoretical guarantee; and the other is the better algorithm convergence and smaller RMSEs than the inflation method as shown in the simulation studies and the empirical study on sea temperature assimilation. The latter means that a smaller ensemble size for the EnKF can be used at a given level of accuracy which is particularly attractive as the running of the geophysical models is generally very expensive. The proposed high dimensional EnKF offers significant potential for application in earth science where the geophysical models, the state variables or the observations are ultra-high dimensional, such as global carbon calculation, ocean data assimilation, weather forecasting, sea ice prediction and the generation of high-resolution data sets.
{acks}
[Acknowledgments] The authors would like to thank Professor Xiaogu Zheng for insightful discussions and comments.
{funding}
The research was partially supported by a National Natural Science Foundation of China grant Nos. 12292980 and 12292983. This research was also partially supported by the National Key Scientific and Technological Infrastructure project “Earth System Science Numerical Simulator Facility” (EarthLab).
[…]

### v2

@@ 7 Discussion
In this paper, we propose HD-EnKF algorithms to suit the needs of high dimensional data assimilation. We take the opportunity to study the theoretical properties of the EnKF under both the fixed and the high dimensional cases for both the correct and misspecified models. The study shows that the proposed HD-EnKF can replace the widely used inflation method for data assimilation of a large dynamic system based on two considerations. One is that the established theoretical analysis on the consistency and the error bounds makes the proposed HD-EnKF methods more understandable with a theoretical guarantee; and the other is the better algorithm convergence and smaller RMSEs than the inflation method as shown in the simulation studies and the empirical study on sea temperature assimilation. The latter means that a smaller ensemble size for the EnKF can be used at a given level of accuracy which is particularly attractive as the running of the geophysical models is generally very expensive. The proposed HD-EnKF has great potential for applications in earth science where the geophysical models, the state variables or the observations are ultra-high dimensional, such as global carbon calculation, ocean data assimilation, weather forecasting, sea ice prediction and the generation of high-resolution data sets.
SUPPLEMENTARY MATERIAL
We provide detailed steps of the proposed HD-EnKF algorithm, bandwidth selection of the mid-banding estimator, additional theoretical results and simulation results, and the details for the proofs of the theoretical results in the Supplementary.